Another question from my CCT part 2 that i did not understand at all.
Three people are standing apart from each other in the frm of a triangle. Fred is 5cm away from Angela 6cm away from Mark and Mark is 7cm away from fred. What angles do they make.

Answer by addingup(3677)

our sides are a, b, and c
our angles are A, B, and C
a = 5
b = 6
c = 7
-----------------------------
cos C = (a^2+b^2-c^2)/(2ab)
cos C = (5^2+6^2-7^2)/(2*5*6) = 0.2
C = cos^-1(0.2) = 78.46 degrees
-------------------
cos A = b^2+c^2-a^2)/(2bc)
cos A = (6^2+7^2-5^2)/(2*6*7) = 0.714
A = cos^-1(0.74) = 42.27 degrees
------------------------
to find our third angle, let's remember that the sum of the angles of EVERY triangle equal 180 degrees:
180-(78.46+42.27) = 59.27 degrees
So there you have it, those are your three angles.